Question 20 1 pts What is the absolute value of the Jacobian of: x = uv, y = uż + v2 at the input point (u, v) = (2,3)?
Answer #1

Given! y = y2 + 2 To find! Absolute value of jacobian at (1:0) = (213) solution: If x= x(uiv) and y=f(u) then its Jacobian is